40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 4 02:13:02 2020, gautam, Update, LSC, Locking updates modifiedOutMat.pngOLTFs.pdf
    Reply  Wed Feb 5 16:35:12 2020, gautam, Update, LSC, Diagnosis plan ALSnoise.pdf
       Reply  Thu Feb 6 01:16:58 2020, gautam, Update, LSC, Diagnosis results IMCspec_ALS.pdfIMCspec_ALS_smallSpan.pdf
    Reply  Thu Feb 6 01:25:50 2020, gautam, Update, LSC, Locking updates semiStableArms.pngarmAngStability.png
       Reply  Fri Feb 7 02:24:24 2020, gautam, Update, LSC, Some short notes 
Message ID: 15192     Entry time: Thu Feb 6 01:25:50 2020     In reply to: 15185     Reply to this: 15195
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking updates 

Summary:

I managed to partially stabilize the arm citculating powers - they stay in a region in which the REFL 11 signal is hopefully approximately linear and so I can now measure some loop TFs and tweak the transition appropriately. 

Procedure:

The main change I made tonight was to look at the REFL11 signal as I swept the ALS CARM offset through 0. I found that the maximum arm powers coincided with a non-zero REFL11 signal value (i.e. a small CARM offset was required at the input to the CARM_B filter bank). Not so long ago, I had measured the PM/AM ratio for 11 MHz to be ~10^5 - so it's not entirely clear to me where this offset is coming from. Then, I was able to turn on the integrator (z:p = 20:0) in the CARM_B filter bank while maintaining high POP_DC. At this point, I ramped up the IN2 gain on the IMC servo board (= AO path), and was able to further stabilize the power. 

Attachment #shows this sequence from earlier in the evening. Note that in this state, both ALS and IR control of CARM is in effect. The circulating power is fluctuating wildly - partly this is probably the noisy ALS control path, but there is also the issue of the (lack of) angular control - although looking at the transmon QPDs and the POP QPD signals, they seem pretty stable.

The next step will be to try and turn off the ALS control path. Eventually, I hope to transition DARM control to AS55 as well. But at this point, I can at least begin to make sense of some of the time series signals, and get some insight into how to improve the lock.

Quote:

No systematic diagnosis scheme comes to mind.

Attachment 1: semiStableArms.png  324 kB  Uploaded Thu Feb 6 18:08:58 2020  | Hide | Hide all
semiStableArms.png
Attachment 2: armAngStability.png  260 kB  Uploaded Thu Feb 6 18:30:45 2020  | Hide | Hide all
armAngStability.png
ELOG V3.1.3-